# Incorporating Judgment Prediction into Legal Case Retrieval via Law-aware Generative Retrieval

### v1

@@ 2.4. Law Structure Constraint Tree
Considering that both legal case retrieval and judgment prediction require guidance by the law where there usually exists a “Chapter-Section-Article” hierarchy, we argue that the decision process to judge whether two cases are relevant in case retrieval is analogous to the search in such tree-like hierarchy. In other words, when the practitioners search for the relevant cases given a query case, they always expect the charge of the relevant precedent case to be located at the same position in the hierarchy as the charges applicable in the query case. Therefore, we devise a module $f_{T}$ that leverages the inherent hierarchy of the law to construct a law structure constraint tree $T$ as shown in Figure 3 and assigns the law-aware semantic identifier $id$ for each $c\in\mathcal{C}$:
@@
where $id$ is in the prefix-suffix style. The prefix depends on the position of the applicable charges within the tree. As illustrated in Figure 3, case 809 involves the charge of “crime of robbery” which falls under Article 269 of Chapter 5 of Section 2 of the Criminal Law of the People’s Republic of China. Hence, the assigned prefix for this case is 0-2-5-269, where 0 represents the root node of the tree. Regarding the suffix, we regard legal cases are the children of their corresponding charges on the tree, and assign a unique identifier to each case under a crime node. Compared to current works that employ hierarchical $k$-means to create identifiers for each document, ours avoid that same integers may have different meanings at different levels, and thus ensure the effectiveness of model learning.
In the retrieval (inference), the identifiers of top relevant cases can be easily generated via beam search. Due to the hierarchical nature of laws as analyzed above, it is convincing to use constrained beam search (Anderson et al., 2017; Hu et al., 2019; Post and Vilar, 2018) decoding process with our law structure constraint tree, which in turn only generates the valid identifiers. On the other hand, one unique feature of the legal domain is that a single case can involve multiple charges. When the query is this kind of case, the ideal retrieval results should encompass these charges. Therefore, we assign $k$ identifiers to cases involving $k$ charges, with each identifier corresponding to a specific charge. For example, as shown in Figure 3, the case 725 involves charge 267 and charge 269, the valid identifiers for this case include 0-2-5-267-725 and 0-2-5-269-725. In doing so, during the retrieval, we expect the model to retrieve all identifiers of the target case, thereby increasing the probability of the target being retrieved.
@@ 2.5. Revision Loss
Besides employing the typical language model training objectives (Equation 1 and Equation 2), we also develop a revision loss for the consistency between the query and retrieved cases, i.e., we aim at directly minimizing their charge-level and case-level discrepancy.
Formally, as illustrated in Figure 3, given the predicted identifier (list of integers) $[\hat{id}_{1},\hat{id}_{2},\cdots,\hat{id}_{L}]$ and the corresponding ground-truth identifier $[{id}_{1},{id}_{2},\cdots,{id}_{L}]$ for a query $q$ at the same position of the ranking, both having a length of $L$, we compare the difference between them and calculate the reward $R_{t}$ at each step $t\in[1,L]$ as follows:
@@
where $\mu$ is the constant reward unit, $\gamma_{h}\in(0,1]$ is the hierarchy penalty factor used to penalize the differences between predictions and labels layer by layer along the law structure constraint tree, with larger penalties as it get closer to the tree root and smaller penalties as it get closer to the tree leaves. Intuitively, if two cases share the same prefix, they are likely relevant to each other because of the same applicable charge, thus receiving higher rewards. Conversely, if two cases belong to different chapters in terms of applicable charges, they are likely dissimilar resulting in lower rewards.
@@
where $\gamma_{r}\in(0,1]$ the optional rank penalty factor used to focus on the top of ranking; $bz$ denotes the beam size, i.e., the number of cases to retrieve for each query; $p(\hat{id}_{t}|q)$ is the probability that predicting to generate $\hat{id}_{t}$ given $q$ at layer $t$; $\lambda_{w}$ is the balance coefficient. To further handling the sparse reward issue and improve the training efficiency, we follow (Xin et al., 2020; Chen et al., 2021; Liu et al., 2023a) and add the second term to Equation 9 that directly increases the probability of generating ${id}_{t}$. Therefore, the overall training objective for Gear is:
@@
2.6. Discussion
In this part, we highlight the merits and limitations of the proposed Gear approach as follows:
Transparency. Inspired by the decision-making process in legal scenarios, our approach formulates the legal case retrieval as the traversal on the law structure constraint tree. This design not only provides valuable contextual information but also improves the transparency of retrieval results by tracing back the decision-making process.
Versatility. Since the retrieved cases are represented by the law-aware semantic identifier with each digit denoting one prediction at each layer in the law structure constraint tree, we can directly predict the applicable judgments for the query case according to the retrieval results. This design makes it versatile to unify legal case retrieval and judgment prediction.
Limitation. In this work, we mainly focus on the Criminal Law of the People’s Republic of China. It follows the civil law system and has an explicit hierarchy. In countries that follow the common law system such as the United States and the United Kingdom, their laws usually do not possess an explicit hierarchical structure. It may limit the applicability of our Gear. Fortunately, for example, in the United States, there is a combination of federal and state laws, where the latter often follows a structure involving “titles, chapters, sub-chapters” hierarchy. Therefore, our method can be applied to such scenarios.
@@ 3. Experiments
In this section, we conduct experiments to answer the following research questions:
  • RQ1: How does Gear perform on legal case retrieval compared to state-of-the-art methods?
  • RQ2: How effective are the three modules in Gear?
  • RQ3: Can Gear show competitive performance on judgment prediction?
  • RQ4: What is the quality of the rationales extracted by Gear including effectiveness and efficiency?
  • RQ5: Can Gear incur less time overhead in legal case retrieval compared to popular generative methods?
@@ 3.1.1. Datasets
We conduct experiments on LeCaRDv2(Ma et al., 2021) and ELAM(Yu et al., 2022) for legal case retrieval. Table 1 reports basic statistics of both datasets. We leave the detailed descriptions in Appendix A.1.
Table 1. Statistics of ELAM and LeCaRDv2. “(Top 1)” indicates the charges corresponding to the most relevant candidate case serve as the label charges for the query case.
@@
Statistics ELAM LeCaRDv2
Avg. length per candidate case 1163.68 1568.38
Avg. length per query case 1304.98 558.18
Avg. # charges per candidate case 1.00 1.50
Avg. # charges per query case (Top 1) 1.00 1.53
# Train candidate cases/query 1332 1390
# Query cases involved 147 653
# Crimes involved 97 100
[…]
@@
ANCE (Xiong et al., 2020) 0.0769 0.1894 0.2357 0.3300 0.1512 0.0362 0.0929 0.0968 0.1145 0.1323
BERT (Devlin et al., 2018)+ANN 0.0302 0.0908 0.1405 0.2461 0.0921 0.0299 0.0718 0.0757 0.0934 0.1162
Lawformer (Xiao et al., 2021a)+ANN 0.0484 0.1638 0.2209 0.3185 0.1410 0.0318 0.0591 0.0728 0.1093 0.1238
SAILER (Li et al., 2023) 0.0670 0.2132 0.3282 0.4604 0.1729 0.0208 0.0844 0.1510 0.2371 0.1318
@@
Gear ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
@@
3.1.2. Baselines
We compare Gear with three types of representative baselines in this work including term-based methods, dense retrieval methods, and generative retrieval methods. We leave the introduction on baselines and the detailed implementation in Appendix A.2 and Appendix A.3, respectively.
@@ 3.1.3. Evaluation Metrics
Following previous works (Tay et al., 2022; Zhuang et al., 2022; Wang et al., 2022; Zhou et al., 2022), we leverage the commonly adopted metrics, including Recall (R) and MRR. The averaged results on all test cases are reported.
To demonstrate Gear’s ability on judgment prediction, we introduce a new metric called $coverage@k$, assessing the percentage of charges involved in the query cases that are covered by top-$k$ retrieved cases. This metric evaluates the charge-level consistency between the query case and the retrieved cases. Considering that query cases are usually undecided and LeCaRDv2 does not have the ground-truth judged charges for query cases, we adopt $coverage_{\text{freq}}$ and $coverage_{\text{top1}}$ to create a pseudo label for each query case. In terms of $coverage_{\text{freq}}$, for each query case, we create a charge dictionary from all the charges involved in all relevant candidate cases and select the charge that appears most frequently as the charge label for the query. In terms of $coverage_{\text{top1}}$, for each query case, we take the charge involved in the most relevant candidate case as the charge label for the query. In our experiments, we compute $coverage@k$ metric at 1, 3, 5, and 10.
@@ 3.2. Retrieval Performance
$\bullet$ Gear demonstrates a significant performance advantage over all baseline methods on both datasets. The relative improvements in terms of MRR on the ELAM and LeCaRDv2 are at least 36.77% and 16.09%, respectively. These results indicate Gear’s effectiveness. We contribute the improvement to the elaborate design for incorporating judgment into case retrieval including rationale extraction, identifier construction, and the revision loss. See Section 3.3 for detailed analysis for each module of Gear.
$\bullet$ Compared to sparse retrieval and dense retrieval methods, current generative retrieval methods struggle to achieve satisfying performance in the legal domain. Generative retrieval baselines merely focus on mapping original cases to their $k$-means style identifiers. Without injecting the law knowledge, these methods inevitably yield sub-optimal results. SAILER uses the case-level structure and obtains the second-best performance in ELAM where both query and candidate cases have clear structure. In LeCaRDv2 where only the fact part is provided for query case and there exhibits a distribution shift between query and candidate, BM25 performs second best. In contrast, Gear explicitly leverages the law knowledge to extract rationales, construct identifiers, and involve judgment prediction for legal case retrieval. These designs lead to significant performance improvement.

### v2

@@ 3.4. Law Structure Constraint Tree
Given that both legal document retrieval and judgment prediction require guidance by law, typically organized in a “Chapter-Section-Article” hierarchy, we argue that the decision process to judge whether two documents are relevant in document retrieval is analogous to the search in such a tree-like hierarchy. In other words, when legal practitioners search for the relevant documents given a query, they always expect the charge of relevant documents to be located at the same position in the law hierarchy as the charges applicable in the query. Therefore, we devise a module $f_{T}$ that leverages the inherent hierarchy of the law to construct a law structure constraint tree $T$ as illustrated in Figure 2 and assigns the law-aware semantic ID $id$ for each $c\in\mathcal{C}$:
@@
where $id$ is in the prefix-suffix style. The prefix depends on the position of the applicable charges within the tree. As shown in Figure 2, document 809 involves the charge of “crime of robbery” which falls under Article 269 of Chapter 5 of Section 2 of the Criminal Law of the People’s Republic of China. Hence, the assigned prefix for this document is 0-2-5-269, where 0 represents the root node of the tree. As for the suffix, we regard documents are the children of their corresponding charges on the tree, and assign a unique ID to each of them under a crime node. Compared to current works that employ hierarchical $k$-means to create IDs for each document, ours avoid same integers may have different meanings at different levels, and thus ensure the effectiveness of model training.
On the other hand, one unique feature of the legal domain is that a single document can involve multiple charges. When the query is this kind of document, the ideal retrieval results should encompass these charges. Therefore, we assign $k$ IDs to documents involving $k$ charges, with each ID corresponding to a specific charge. For example, as shown in Figure 2, the 725 involves charge 267 and charge 269, the valid IDs for this document include 0-2-5-267-725 and 0-2-5-269-725. In doing so, during retrieval, we expect the model to retrieve all IDs of the target documents, thereby increasing the probability of the target being retrieved.
@@ 3.5. Revision Loss
Besides employing the typical language model training objectives (Equation 1 and Equation 2), we also develop a novel training objective called the revision loss for the consistency between the query and retrieved documents, i.e., we aim to directly minimize their judgment-level and document-level discrepancy.
Formally, as illustrated in Figure 2, given the predicted ID (list of integers) $[\hat{id}_{1},\hat{id}_{2},\cdots,\hat{id}_{L}]$ and the corresponding ground-truth ID $[{id}_{1},{id}_{2},\cdots,{id}_{L}]$ for a query $q$, both having a length of $L$, we compare the difference between them and calculate the reward $R_{t}$ at each step $t\in[1,L]$ as follows:
@@
where $\mu$ is the constant reward unit, $\gamma_{h}\in(0,1]$ is the hierarchy penalty factor used to penalize the differences between predictions and labels layer by layer along the law structure constraint tree, with larger penalties as it gets closer to the tree root and smaller penalties as it gets closer to the tree leaves. Intuitively, if two documents share the same prefix, they are likely relevant to each other because of the same applicable charge, receiving higher rewards.
@@
where $\gamma_{r}\in(0,1]$ the optional penalty factor used to focus on the top of retrieved document list; $bz$ denotes the beam size, i.e., the number of documents to retrieve for each query; $p(\hat{id}_{t}|q)$ is the probability that predicting to generate $\hat{id}_{t}$ given $q$ at layer $t$; $\lambda_{w}$ is the balance coefficient. To further handle the sparse reward issue and improve the training efficiency, we follow (Xin et al., 2020; Chen et al., 2021; Liu et al., 2023a) and add the second term to Equation 9 that directly increases the probability of generating ${id}_{t}$. Thus, the overall training objective is:
@@
3.6. Inference
In the inference, we aim to retrieve the top-$k$ documents from the candidate pool. Since we have assigned hierarchical semantic ID to each document based on the law-aware constraint tree where each leaf node corresponds to a candidate document in the pool, we utilize the constrained beam search (Anderson et al., 2017; Hu et al., 2019; Post and Vilar, 2018) to ensure all the generated document IDs are valid within the tree.
@@ 4. Experiments
In this section, we conduct experiments to answer the following research questions: RQ1: How does GEAR perform on legal document retrieval compared to state-of-the-art methods? RQ2: How effective are the three modules in GEAR? RQ3: Can GEAR show competitive performance on judgment (applicable charges) prediction? RQ4: What is the quality of the rationales extracted by GEAR including effectiveness and efficiency? RQ5: Can GEAR incur less time overhead in legal document retrieval compared to popular generative methods? RQ6: How robust is GEAR across languages and domains (e.g. in statutory article retrieval)?
@@ 4.1.1. Datasets
ELAM [^3] https://github.com/ruc-wjyu/OPT-Match. (Yu et al., 2022b) is a Chinese LCR dataset, focusing on criminal cases. ELAM has corresponding labels for both case retrieval and judgment prediction, which is suitable for our goal. We exclude those cases with multiple applicable charges to consider the retrieval and judgment prediction performance in a single charge scenario. The resulting candidate pool size of ELAM is 1332. Other data preprocessing is aligned to (Li et al., 2023a).
LeCaRDv2444https://github.com/THUIR/LeCaRDv2. (Li et al., 2023b) is the official updated version of LeCaRD  (Ma et al., 2021). In this dataset, the relevance labels are divided into four levels, ranging from 3 to 0, indicating a gradual decrease in relevance. We follow the data preprocessing approach of (Li et al., 2023a), with the exception of increasing the candidate pool size from 100 to 1390 to further validate the effectiveness of baselines and our model. In LeCaRDv2, cases encompass both single and multiple charges, averaging 1.5 charges per case. Considering the ground-truth judgment labels of query cases have not been provided, we ask two legal experts (Ph.D. in Law) to annotate the charge label for the testing queries. The experts are proficient in Chinese criminal law with sufficient experience in handling cases similar to this dataset. They carefully align the LeCaRDv2 judgment criteria (Li et al., 2023b) before annotation and discuss opinions to reach a consensus, ensuring accurate labeling.
BSARD [^5] https://huggingface.co/datasets/maastrichtlawtech/bsard. (Louis and Spanakis, 2021) is a SAR dataset composed of more than 1.1K legal questions labeled by domain experts with relevant articles selected from the 22K law articles gathered from 32 publicly available Belgian codes. It is worth noting that BSARD contains structural annotations of corresponding laws, facilitating the utilization of law structural knowledge.
4.1.2. Baselines
We consider three types of baselines in this study.
(1) Sparse retrieval methods: Query Likelihood (QL) (Zhai et al., 2008) is a probabilistic language modeling approach employed to assess the relevance of documents to a provided query. BM25 (Robertson et al., 2009) is a probabilistic information retrieval model widely used in the field of text retrieval. BM25 takes into account both term frequency and document length normalization.
(2) Dense retrieval methods: BERT (Devlin et al., 2018) is a strong baseline in ad-hoc retrieval tasks in the open domain. In this paper, we adopt the checkpoint that is pre-trained on a large Chinese corpus Following (Tay et al., 2022; Wang et al., 2022; Zhuang et al., 2022), after encoding legal documents using BERT, we then apply Approximate Nearest Neighbor (ANN) search algorithms to retrieve relevant documents. Legal-BERT [^6] https://github.com/thunlp/OpenCLaP. (Chalkidis et al., 2020) is a variant of BERT that undergoes specific training in the legal domain to better understand and process text related to law. Lawformer  (Xiao et al., 2021a) is a Longformer (Beltagy et al., 2020) backbone pre-trained on large legal case corpus, to encode legal texts. ChatLaw-Text2Vec [^7] https://huggingface.co/chestnutlzj/ChatLaw-Text2Vec. (Cui et al., 2023) is a legal text matching model based on ChatLaw which is pre-trained on a corpus of 936,727 legal documents. G-DSR (Louis et al., 2023) uses legal-CamemBERT [^8] https://huggingface.co/maastrichtlawtech/legal-camembert-base., a legal variant of CamemBERT trained on BSARD dataset. It takes into account both the dense representation of text and the graph representation of legal structures. G-DSR is the state-of-the-art SAR method in the French legal domain. SAILER  (Li et al., 2023a) is a structure-aware LCR model. It adopts an asymmetric encoder-decoder architecture to integrate structures of legal case document information into dense vectors. SAILER achieves state-of-the-art retrieval performance in Chinese LCR domain. As for the training of all baselines, we follow (Tay et al., 2022; Wang et al., 2022; Zhuang et al., 2022) and use Approximate nearest neighbor Negative Contrastive Estimation (ANCE) (Xiong et al., 2020) method.
(3) Generative retrieval methods: DSI (Tay et al., 2022) is a new paradigm for document retrieval tasks. It utilizes a Transformer-based encoder-decoder model to map queries directly to relevant IDs. DSI achieves the end-to-end retrieval. NCI (Wang et al., 2022) improves DSI in terms of using constrained beam search and prefix-aware weight-adaptive decoder. Both DSI and NCI use hierarchical $k$-means clustering of document vectors to create $k$-means IDs. DSI-QG (Zhuang et al., 2022) design a query generation process on the top of DSI, which can mitigate data distribution mismatches present between the indexing and the retrieval phases. Ultron (Zhou et al., 2022) improves DSI through adopting product quantization to create semantic IDs and using URLs to create term-based IDs.
4.1.3. Implementation Details
We implement baseline methods following the suggestions in the original papers.
(1) For classical term-based baselines, we use the pyserini and genism toolkits with the default parameters.
(2) For dense retrieval baselines, we use the open-sourced checkpoint to initialize model parameters of the pre-trained models and use faiss toolkit to implement ANN algorithms. The batch size is set to 16. The max length of the input text is set to 1024 for Lawformer and 512 for the other models. We tune these models with in-batch contrastive loss.
(3) For generative retrieval baselines, we directly use their official open-source implementations, employing the pre-trained T5 “Randeng” [^9] https://huggingface.co/IDEA-CCNL/Randeng-T5-77M-MultiTask-Chinese. as backbone for Chinese legal domain, and “t5-base” [^10] https://huggingface.co/t5-base. Please note that this is applicable to French datasets. for the French legal domain. We use beam search to retrieve relevant cases, where the beam size is set to 30.
We keep the backbone model and beam size same as baselines and set the max input length of GEAR to 512, the rest hyperparameters are tuned as follows: the batch size is set to 2; the learning rate is tuned from [1e-5,1e-4] with step size 2e-5; in the rationale extraction module, $k_{1}$, $k_{2}$, $k_{3}$ are respectively set to 2, 5, 15 for ELAM and 10, 20, 15 for LeCaRDv2; $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ are respectively set to 10.0, 1.0, 0.1; in the law structure constraint tree module, the height of the tree (the length of the hierarchical ID) $L$ is set to 4; in the training, $\mu$ is set to 1; $\gamma_{h}$ is tuned from $\{0.01,0.1\}$; $\gamma_{h}$ is set to 1; $\lambda_{w}$ is tuned from $\{1,10,100\}$; $\lambda_{l}$ is tuned from [1e-4,1e-2] with step size 5e-4. Hyperparameters of GEAR are tuned using grid search with Adam(Kingma and Ba, 2014). ALL experiments are conducted on a single NVIDIA RTX A6000. The source code and datasets have been shared at: https://github.com/E-qin/GEAR.
Table 1. Statistics of the datasets. Since queries in BSARD do not involve judgments, we omit the judgment prediction comparison on this dataset.
@@ 4.1.3. Implementation Details
Statistics ELAM LeCaRDv2 BSARD
Avg. length per candidate document 1163.68 1568.38 880.29
Avg. length per query document 1304.98 558.18 92.48
Avg. # charges per candidate case 1.00 1.50 -
# Available candidates per query 1332 1390 1612
# Query documents involved 147 653 1108
# Charges involved of judgment 97 100 -
Table 2. Performance comparisons of our approach and the baselines on ELAM dataset and LeCaRDv2 dataset. The best and the second-best performances are denoted in bold and underlined fonts, respectively. “R@K” is short for “Recall@K”. denotes GEAR performs signifcantly better than baselines based on two-tailed paired t-test with Bonferroni correction ($p<0.05$).
@@
BERT (Devlin et al., 2018) 0.0302 0.1008 0.1405 0.2861 0.1521 0.0299 0.0718 0.1557 0.2534 0.1162
Legal-BERT (Zhong et al., 2019) 0.0384 0.0938 0.1509 0.3123 0.1510 0.0218 0.0620 0.1081 0.2743 0.1138
Lawformer (Xiao et al., 2021a) 0.0537 0.1682 0.2220 0.3789 0.1701 0.0518 0.1491 0.2728 0.3593 0.1638
ChatLaw-Text2Vec (Cui et al., 2023) 0.0385 0.1371 0.2065 0.3323 0.1694 0.0356 0.0813 0.1510 0.3380 0.1379
SAILER (Li et al., 2023a) 0.0729 0.2132 0.3282 0.4604 0.2029 0.0608 0.1644 0.2910 0.4271 0.2018
@@
GEAR 0.0793 0.2368 0.3356 0.4976 0.2365 0.0630 0.1706 0.3142 0.4625 0.2162
@@ 4.1.4. Evaluation Metrics
For a fair comparison, we follow previous works (Tay et al., 2022; Zhuang et al., 2022; Wang et al., 2022; Zhou et al., 2022) and leverage the commonly adopted metrics, including Recall (R) and MRR. The averaged results on all test cases are reported.
To demonstrate GEAR’s ability on judgment prediction, we introduce a new metric called $coverage@k$, assessing the percentage of charges involved in the query that are covered by top-$k$ retrieved documents. This metric evaluates the charge-level consistency, i.e., the extent to which the retrieval process contributes to the efficacy of legal judgment. Formally, $coverage@k$ is defined as:
(11) $coverage@k=\frac{1}{N_{q}}\sum_{i=1}^{N_{q}}\frac{|\mathcal{E}_{i}^{\text{Top-}k}\cap\mathcal{E}_{i}|}{|\mathcal{E}_{i}|},$
where |||~{}|| | is the size of a set; for $i$-th testing query, $\mathcal{E}_{i}$ denotes its label set of applicable judgment charges, $\mathcal{E}_{i}^{\text{Top-}k}$ denotes the the set of charges of the top-$k$ retrieved documents; $N_{q}$ is the number of testing queries. In our experiments, we compute $coverage@k$ metric at 1, 3, 5, and 10.
@@ 4.2. Retrieval Performance
(1) GEAR demonstrates a significant performance advantage over all baseline methods on both datasets. The relative improvements of MRR on the ELAM and LeCaRDv2 are at least 16.55% and 7.13%, respectively. These results indicate GEAR’s effectiveness. We attribute the improvement to the elaborate design for explicitly integrating judgment into case retrieval including rationale extraction, law-aware ID assignment, and the revision loss. See Section 4.3 for the detailed analysis of each module of GEAR. (2) Compared to sparse retrieval and dense retrieval methods, current generative retrieval methods struggle to achieve satisfying performance in the legal domain. Without injecting the law knowledge, generative retrieval baselines inevitably yield sub-optimal results. On the other hand, SAILER introduces an implicit training objective that uses the fact description of the legal document to predict its judgments. In this way, SAILER produces judgment-aware document representations and achieves the second-best performance. However, SAILER fails to provide explicit evidence of judgment consistency for relevance modeling, leading to an inferior retrieval performance compared to GEAR.
(a) $coverage$ on ELAM
(b) $coverage$ on LeCaRDv2